OpenStudy (anonymous):

Assume that in the circuit above, V0=10V, T=0.5ms, Rw=1Ω, Lw=0.5mH and RL=4Ω

OpenStudy (anonymous):

anyone??

OpenStudy (anonymous):

@salman.abid12 , just a sec, I'll calculate for you

OpenStudy (anonymous):

@salman.abid12 C - 7.94645719261

OpenStudy (anonymous):

@salman.abid12 D - 0.05354280739

OpenStudy (anonymous):

sure???

OpenStudy (anonymous):

Dude, I'm working on this in the airport :)) I'm the guy that posted the formulas here (but they were deleted), so the formulas are correct. Hope my calculations are correct too, but I don't have a calculator.

OpenStudy (anonymous):

V0=5V, T=0.5ms, Rw=1Ω, Lw=1mH and RL=1Ω. ???

OpenStudy (anonymous):

@rmpk means there is some risk... ok let me chek it..

OpenStudy (anonymous):

yupeeeeee! thanks a lot dude @rmpk..... that were correct.

OpenStudy (anonymous):

No problem, glad it was fine.

OpenStudy (anonymous):

Q1 a)ia=2*i1*R/(2*R+rm) and vo=ia*rm b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm c) part a) + part b) d)2R Q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) Q3 same as HP8p1 Q4 p1=V*R2/(R1+R2)*(1-e^(-(R1+R2)*t/(R1*R2*C))) p2=(V*R2/(R1+R2))*cos(1/sqrt(L*C)*(t-T)) p3=C*(V*R2/(R1+R2))*sin((t-T)*1/sqrt(L*C))/sqrt(L*C) Q5 E V B Y F Z Q6

OpenStudy (anonymous):

What about Q6 A and B?